# Coupled-Cluster Calculations of Infinite Nuclear Matter in the Complete Basis Limit Using Bayesian Machine Learning

### v2

@@ Abstract
Background Infinite nuclear matter provides valuable insights into the behavior of nuclear systems and aids our understanding of atomic nuclei and large-scale stellar objects such as neutron stars. However, partly due to the large basis needed to converge the system’s binding energy, size-extensive methods such as coupled-cluster theory struggle with long computational run times, even using the nation’s largest high-performance computing facilities.
@@ I Introduction
nuclear matter can be studied using different methods. We will use an ab initio many-body method called coupled-cluster theory (CC) Coester (1958); Coester and Kümmel (1960); Kümmel et al. (1978); Bishop (1991); Hagen et al. (2014); Heisenberg and Mihaila (1999); Dean and Hjorth-Jensen (2004); Bartlett and Musiał (2007) in this work. Coupled-cluster theory is a powerful many-body method popular in, among others, quantum chemistry, nuclear physics, and condensed matter physics Taube and Bartlett (2008); Kucharski and Bartlett (1998); Bartlett and Musiał (2007); Bartlett et al. (1990); Raghavachari et al. (1989); Kutzelnigg (1991); Čižek and Paldus (1971); He et al. (2001); Margraf and Reuter (2018); Weiler et al. (2022); Bishop and Lührmann (1981); Mihm et al. (2021). It provides a systematic framework for calculating correlations in addition to the mean field, with typical correlations like one-particle-one-hole (single excitations or just singles), two-particle-two-hole (doubles), three-particle-three-hole (triples) or more complicated excitations that are summed to infinite order. Here we use the coupled-cluster approximation called coupled-cluster doubles with perturbative triples CCD(T) to calculate the binding energy of infinite nuclear matter Hagen et al. (2014); Raghavachari et al. (2013); Shavitt and Bartlett (2009); Urban et al. (1985) [^1] For infinite matter calculations, one-particle-one-hole excitations (singles) are not present.. This approximation, which represents state-of-the art infinite matter calculations with CC theory, calculates the correlation energy using two-particle two-hole excitations exactly while including the effects of three-particle three-hole excitations perturbatively. However, though CCD(T) calculations provide the most accurate results, the calculations have a run time that scales as $O(M^{6})$ for the iterative part and $O(M^{7})$ for the non-iterative part, where $M$ is the number of single-particle states in the calculation. In infinite nuclear matter calculations, the number of single-particle states $M$, which in principle is infinitely large, is typically of the order $M\sim 10^{3}$ or larger in actual implementations. These large single-particle basis sets which in turn define the number of possible many-particle excitations, are needed in order to obtain converged total energies with respect to the number of single-particle states. Due to the large number of intermediate single-particle states needed for accurate CCD(T) calculations, infinite nuclear matter require many individual calculations. Large-scale studies of infinite nuclear matter, spanning a wide range of densities and proton fractions, have prohibitively expensive computational costs He et al. (2001).
@@ II Background
The system of interest in this work, infinite nuclear matter, is an infinite matter system made of nucleons (protons and neutrons) in various fractions which interact via the nuclear force Baardsen et al. (2013). Studies of infinite nuclear matter from the many-body perspective are important for understanding the matter within dense astronomical objects such as neutron stars, which offer insights into nuclear processes and astrophysical observables. However, neutrons stars also contain matter that spans several orders of magnitude ($0.1$fm-3 and greater) and contain neutrons, protons, electrons, and muons (among other particles) in various fractions. Lietz et al. (2017); Ref (1983); Lattimer and Prakash (2001, 2007); Steiner et al. (2010); Lattimer (2012); Heiselberg and Hjorth-Jensen (2000); Drischler et al. (2021a, b); Burgio et al. (2021). These particles exist in beta equilibrium ($\beta$-equilibrium) governed by the weak force. For this work, we will restrict our calculations to around nuclear saturation density (0.16 $fm^{-3}$).
@@
For this work, we will limit our calculations to an infinite matter system containing only protons and neutrons in various ratios. The density of protons in the matter, $\rho_{p}$, compared to the total density, $\rho$, defines the proton fraction, $x_{p}$. Note that $\rho=\rho_{p}+\rho_{n}$, where $\rho_{n}$ is the density of the neutrons is the matter Lietz et al. (2017)
@@
Changing the proton fraction of the infinite nuclear matter system changes its properties. The two types of infinite nuclear matter we will be studying in this work are pure neutron matter (PNM) which has $x_{p}=0$ and symmetric nuclear matter (SNM), which has $x_{p}=0.5$.
From the proton fraction, we can define the symmetry energy as the difference between the energy for pure neutron matter and symmetric nuclear matter at a set density Lietz et al. (2017):
$E_{sym}(\rho)=E(\rho,x_{p}=0)-E(\rho,x_{p}=1/2).$ (2)
In this paper, we are interested in calculating the energy of both pure neutron matter and symmetric nuclear matter, as well as the symmetry energy. The nuclear interaction used in all calculations in this paper is NNLOGO Ekström et al. (2018); Jiang et al. (2020). This potential is an optimized $\Delta$-full interaction that has been calibrated with nuclear matter properties using 17 low-energy coefficients (LECs) to parameterize the interaction at the next-to-next-to leading order (NNLO) level.
@@
The operator $\hat{T}$ is known as the cluster or the correlation operator, and it is the sum of N i-particle i-hole excitation operators, where N is the number of particles in the system:
@@
If Eq. (4) is used in its entirety in a coupled-cluster calculation, then the result of the calculation is equivalent to the result from solving the Schrödinger equation for the system. However, the cluster operator can rarely be entirely used due to the immense computational costs of the higher-order excitation operators. Fortunately, coupled-cluster theory has a simple and physically motivated truncation scheme in which excitation operators over a set level are forced to zero. Thus, we develop the coupled-cluster singles method when $\hat{T}\approx\hat{T}_{1}$, the coupled-cluster singles and doubles method when $\hat{T}\approx\hat{T}_{1}+\hat{T}_{2}$, and the coupled-cluster singles, doubles, and triples method when $\hat{T}\approx\hat{T}_{1}+\hat{T}_{2}+\hat{T}_{3}$ He et al. (2001); Urban et al. (1985).
@@
Even though we are required to truncate the cluster operator due to computational limitations, these truncated cluster operators can still have very long run times. A CCSD calculation has an expected run time of $O(M^{6})$, where M is the number of single-particle states in the calculation, but a CCSDT calculation has an expected run time of $O(M^{8})$. In nuclear many-body calculations, with complex interactions and large single-particle bases, CCSDT calculations are rarely computationally feasible. However, since many aspects of the nuclear interaction emerge at the three-body order or higher, we still want to include aspects of the $\hat{T}_{3}$ operator in our calculations, but with run times more similar to CCSD calculations. One method that can be used to bridge the gap between CCSD and CCSDT is called coupled-cluster singles and doubles with perturbative triples, CCSD(T). In a CCSD(T) calculation, some aspects of the $\hat{T}_{3}$ operator are approximated with the $\hat{T}_{2}$ operator, but the run time is an iterative $O(M^{6})$ (computing the $\hat{T_{1}}$ and $\hat{T}_{2}$ operators) and a non-iterative component with a run time of $O(M^{7})$ (approximating $\hat{T}_{3}$). While these run times are higher than a CCSD calculation, they are significantly lower than the run times for CCSDT calculations, and thus, CCSD(T) is the most accurate and accessible coupled-cluster method commonly used in nuclear applications. Note that, for infinite matter systems, the $\hat{T}_{1}$ operator has no contribution due to symmetry in the momentum of the system, so the CCSD(T) method simply becomes CCD(T).
@@
Even with the lessened run time of a CCD(T) calculation, it is still computationally difficult to perform calculations at a high enough number of single-particle states that the energy of the system converges (the complete basis limit). Figure 1 shows the run time for CCD(T) calculations as a function of the number of single-particle states for both pure neutron matter and symmetric nuclear matter at N = 66 neutrons (132 nucleons) and $\rho=0.16fm^{-3}$. These calculations were run with an Apple Mac Studio using 3.5 GHz M2 Ultra processor. The coupled-cluster code was parallelized and each calculation was performed with 20 compute nodes. We report run times in units of node hour, defined as the run time of the calculation (in hours) multiplied by the number of nodes used in the calculation. This was done to better reflect the true computational time needed by these calculations. As shown in Fig. 1, as the number of single-particle states in the calculation increases, the run time of the calculation drastically increases as well. However, as we can see in Fig. 1, as the number of single-particle states in the calculations is increased the accuracy of the calculation (compared to results in the complete basis limit) also increases. Therefore, if we want high accuracy in our calculations we need to perform calculations at a high number of single-particle states.
@@ III Methodology
As is shown in Fig. 1, as the number of single particle states in the calculation increases, so does the accuracy of the calculation. Phrased another way, as the number of single particle states in the calculation increases, the CCD(T) energy (or correlation energy) converges to a constant value. This is also a property of another many-body method called many-body perturbation theory (MBPT). When performing a CCD(T) calculation, all of the diagrams needed to find the MBPT energy to the second order (MBPT2) are computed as well, thus the MBPT2 energy (or correlation energy) can be found from a CCD(T) calculation with no additional computational time. Since we know that both of these correlation energies (CCD(T) and MBPT2) converge at some large number of single particle states, then at that value of $M$, the ratio of the two correlation energies is a constant which we will refer to as $m$
@@
The constant $m$ is the value we want to predict with our SRE algorithm, since given this value and the MBPT2 calculation at a high value of $M$, we can predict the CCD(T) correlation energy in the complete basis limit. Therefore, the training data we need to generate for our model are calculations for the CCD(T) and MBPT2 correlation energies at a set density and number of particles, but at increasing numbers of single particle states. For the pure neutron matter results, we used three training points per density, with the number of single particle states ranging from 186 to 294. For the symmetric nuclear matter calculations, we used six training points per calculations, with the number of single particle states ranging from 324 to 684. The training data consisted of a CCD(T) calculation and a MBPT2 calculation at the same number of single particle states, divided by each other to make the ratio in Eq. (5) and arranged in sequential and ascending order
@@ V Conclusion
All code and data needed to recreate the plots and calcula- 588 tions in this paper are available at www.github.com/butler-julie/ INM-SRE.

### v3

@@ Abstract
Background Infinite nuclear matter provides valuable insights into the behavior of nuclear systems and aids our understanding of atomic nuclei and large-scale stellar objects such as neutron stars. However, partly due to the large basis needed to converge the system’s binding energy, size extensive methods such as coupled-cluster theory struggle with long computational run times, even using the nation’s largest high-performance computing facilities.
@@ I Introduction
Infinite nuclear matter can be studied using different methods. We will use an ab initio many-body method called coupled-cluster theory (CC) in this work Coester (1958); Coester and Kümmel (1960); Kümmel et al. (1978); Bishop (1991); Hagen et al. (2014); Heisenberg and Mihaila (1999); Dean and Hjorth-Jensen (2004); Bartlett and Musiał (2007). Coupled-cluster theory is a powerful many-body method popular in, among others, quantum chemistry, nuclear physics, and condensed matter physics Taube and Bartlett (2008); Kucharski and Bartlett (1998); Bartlett and Musiał (2007); Bartlett et al. (1990); Raghavachari et al. (1989); Kutzelnigg (1991); Čižek and Paldus (1971); He et al. (2001); Margraf and Reuter (2018); Weiler et al. (2022); Bishop and Lührmann (1981); Mihm et al. (2021). It provides a systematic framework for calculating correlations in addition to the mean field, with typical correlations like one-particle one-hole (single excitations or just singles), two-particle-two-hole (doubles), three-particle-three-hole (triples) or more complicated excitations that are summed to infinite order. Coupled-cluster theory, as is also the case with other post Hartree-Fock methods like many-body perturbation theory Drischler et al. (2019, 2021a), Green’s function theory Carbone et al. (2013); Barbieri and Carbone (2017), in-medium Similarity Renormalization Group theory Hergert et al. (2017) and other approaches Marino et al. (2024), allows for a systematic inclusion of higher-order correlations not included in a mean-field approach. Here we use the coupled-cluster approximation called coupled-cluster doubles with perturbative triples CCD(T) to calculate the binding energy of infinite nuclear matter Hagen et al. (2014); Raghavachari et al. (2013); Shavitt and Bartlett (2009); Urban et al. (1985) [^1] For infinite matter calculations, one-particle-one-hole excitations (singles) are not present.. This approximation, which represents state-of-the art infinite matter calculations with CC theory, calculates the correlation energy using two-particle two-hole excitations exactly while including the effects of three-particle three-hole excitations perturbatively. However, though CCD(T) calculations provide the most accurate results, the calculations have a run time that scales as $O(M^{6})$ for the iterative part and $O(M^{7})$ for the non-iterative part, where $M$ is the number of single-particle states in the calculation. In infinite nuclear matter calculations, the number of single-particle states $M$, which in principle is infinitely large, is typically of the order $M\sim 10^{3}$ or larger in actual implementations. These large single-particle basis sets which in turn define the number of possible many-particle excitations, are needed in order to obtain converged total energies with respect to the number of single-particle states. Due to the large number of intermediate single-particle states needed for accurate CCD(T) calculations, infinite nuclear matter require many individual calculations. Large-scale studies of infinite nuclear matter, spanning a wide range of densities and proton fractions, have prohibitively expensive computational costs He et al. (2001).
@@ II Background
The system of interest in this work, infinite nuclear matter, is an infinite matter system made of nucleons (protons and neutrons) in various fractions which interact via the nuclear force Baardsen et al. (2013). Studies of infinite nuclear matter from the many-body perspective are important for understanding the matter within dense astronomical objects such as neutron stars, which offer insights into nuclear processes and astrophysical observables. However, neutrons stars also contain matter that spans several orders of magnitude ($0.1$fm-3 and greater) and contain neutrons, protons, electrons, and muons (among other particles) in various fractions. Lietz et al. (2017); Ref (1983); Lattimer and Prakash (2001, 2007); Steiner et al. (2010); Lattimer (2012); Heiselberg and Hjorth-Jensen (2000); Drischler et al. (2021b, c); Burgio et al. (2021). These particles exist in beta equilibrium ($\beta$-equilibrium) governed by the weak force. For this work, we will restrict our calculations to around nuclear saturation density (0.16 $fm^{-3}$).
@@
For this work, we will limit our calculations to an infinite matter system containing only protons and neutrons in various ratios. The density of protons in the matter, $\rho_{p}$, compared to the total density, $\rho$, defines the proton fraction, $x_{p}$. Note that $\rho=\rho_{p}+\rho_{n}$, where $\rho_{n}$ is the density of the neutrons is the matter Lietz et al. (2017).
@@
Changing the proton fraction of the infinite nuclear matter system changes its properties. The two types of infinite nuclear matter that we will study in this work are pure neutron matter (PNM), which has $x_{p}=0$ and symmetric nuclear matter (SNM), which has $x_{p}=0.5$.
From the proton fraction, we define here the symmetry energy as the difference between the energy for pure neutron matter and symmetric nuclear matter at a set density Lietz et al. (2017):
$E_{sym}(\rho)=E(\rho,x_{p}=1/2)-E(\rho,x_{p}=0).$ (2)
For an in-depth discussion of the symmetry energy, see, for example, Refs. Baldo and Burgio (2016) and Lattimer (2014). In this paper, we are interested in calculating the energy of both pure neutron matter and symmetric nuclear matter, as well as the symmetry energy. The nuclear interaction used in all calculations in this paper is NNLOGO Ekström et al. (2018); Jiang et al. (2020). This potential is an optimized $\Delta$-full interaction that has been calibrated with nuclear matter properties using 17 low-energy coefficients (LECs) to parameterize the interaction at the next-to-next-to leading order (NNLO) level.
@@
The operator $\hat{T}$ is known as the cluster or correlation operator, and is the sum of the $N$ i-particle i-hole excitation operators, where N is the number of particles in the system:
@@
If Eq. (4) is used in its entirety in a coupled-cluster calculation, then the result of the calculation is equivalent to the result from solving the Schrödinger equation for the system. However, the cluster operator can rarely be entirely used due to the immense computational costs of higher-order excitation operators. Fortunately, coupled-cluster theory has a simple and physically motivated truncation scheme in which excitation operators over a set level are forced to zero. Thus, we develop the coupled-cluster singles method when $\hat{T}\approx\hat{T}_{1}$, the coupled-cluster singles and doubles method when $\hat{T}\approx\hat{T}_{1}+\hat{T}_{2}$, and the coupled-cluster singles, doubles, and triples method when $\hat{T}\approx\hat{T}_{1}+\hat{T}_{2}+\hat{T}_{3}$ He et al. (2001); Urban et al. (1985). With the inclusion of triples correlations, coupled-cluster theory allows for a computationally efficient inclusion of important many-body correlations beyond a mean-field approach.
@@
Even though we are required to truncate the cluster operator due to computational limitations, these truncated cluster operators can still have very long run times. A CCSD calculation has an expected run time of $O(M^{6})$, where M is the number of single-particle states in the calculation, but a CCSDT calculation has an expected run time of $O(M^{8})$. In nuclear many-body calculations, with complex interactions and large single-particle bases, CCSDT calculations are rarely computationally feasible. However, since many aspects of the nuclear interaction emerge at the three-body order or higher, we still want to include aspects of the $\hat{T}_{3}$ operator in our calculations, but with run times more similar to CCSD calculations. One method that can be used to bridge the gap between CCSD and CCSDT is called coupled-cluster singles and doubles with perturbative triples, CCSD(T). In a CCSD(T) calculation, some aspects of the $\hat{T}_{3}$ operator are approximated with the $\hat{T}_{2}$ operator, but the run time is an iterative $O(M^{6})$ (computing the $\hat{T_{1}}$ and $\hat{T}_{2}$ operators) and a non-iterative component with a run time of $O(M^{7})$ (approximating $\hat{T}_{3}$). Although these run times are higher than for a CCSD calculation, they are significantly lower than the run times for CCSDT calculations, and thus CCSD(T) is the most accurate and accessible coupled-cluster method commonly used in nuclear applications. Note that, for infinite matter systems, the $\hat{T}_{1}$ operator has no contribution due to symmetry in the momentum of the system, so the CCSD(T) method simply becomes CCD(T).
@@
Even with the lessened run time of a CCD(T) calculation, it is still computationally difficult to perform calculations at a high enough number of single-particle states that the energy of the system converges (the complete basis limit). Figure 1 shows the run time for CCD(T) calculations as a function of the number of single-particle states for both pure neutron matter and symmetric nuclear matter at N = 66 neutrons (132 nucleons) and $\rho=0.16fm^{-3}$. These calculations were run with an Apple Mac Studio using 3.5 GHz M2 Ultra processor. The coupled-cluster code was parallelized and each calculation was performed with 20 compute nodes. We report run times in units of node hours, defined as the run time of the calculation (in hours) multiplied by the number of nodes used in the calculation. This was done to better reflect the true computational time needed by these calculations. As shown in Fig. 1, as the number of single-particle states in the calculation increases, the run time of the calculation also drastically increases. However, as we can see in Fig. 1, as the number of single-particle states in the calculations is increased, the accuracy of the calculation (compared to results in the complete basis limit) also increases. Therefore, if we want high accuracy in our calculations we need to perform calculations at a high number of single-particle states.
@@ III Methodology
As is shown in Fig. 1, as the number of single particle states in the calculation increases, so does the accuracy of the calculation. Phrased another way, as the number of single particle states in the calculation increases, the CCD(T) energy (or correlation energy) converges to a constant value. This is also a property of another many-body method called many-body perturbation theory (MBPT). When performing a CCD(T) calculation, all of the diagrams needed to find the MBPT energy to the second order (MBPT2) are computed as well, thus the MBPT2 energy (or correlation energy) can be found from a CCD(T) calculation with no additional computational time. We can use the MBPT2 correlation energy as an estimate of the system’s correlation energy. Even though MBPT2 calculations of infinite nuclear matter can be computationally expensive, they are not nearly as expensive as CCD(T) calculations of the same system and provide a reliable estimate of the intermediate states. Thus, we aim to develop a method which can predict the CCD(T) energy of a system in the complete basis limit, using only MBPT2 calculations and CCD(T) calculations at smaller basis sizes.
Since we know that both of these correlation energies (CCD(T) and MBPT2) converge at some large number of single particle states, then at that value of $M$, the ratio of the two correlation energies is a constant which we will refer to as $m$.
@@
The constant $m$ is the value we want to predict with our SRE algorithm, since given this value and the MBPT2 calculation at a high value of $M$, we can predict the CCD(T) correlation energy in the complete basis limit. Therefore, the training data we need to generate for our model are calculations for the CCD(T) and MBPT2 correlation energies at a set density and number of particles, but at increasing numbers of single particle states. For the pure neutron matter results, we used three training points per density, with the number of single particle states ranging from 186 to 294. For the symmetric nuclear matter calculations, we used six training points per calculations, with the number of single particle states ranging from 324 to 684. The training data consisted of a CCD(T) calculation and a MBPT2 calculation at the same number of single particle states, divided by each other to make the ratio in Eq. (5) and arranged in sequential and ascending order.
@@ V Conclusion
All code and data needed to recreate the plots and calculations in this paper are available at www.github.com/butler-julie/ INM-SRE.